OpenStudy (anonymous):

Solve the exponential equation. 1/16 = 644x – 3

OpenStudy (anonymous):

1/16=64^4x-3

OpenStudy (anonymous):

1/16+3=64^4x (1/16+3)/(64^4)=x

OpenStudy (anonymous):

i got a decimal?

OpenStudy (anonymous):

@bazilham

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

0.0000001825392246246337890625

OpenStudy (anonymous):

1/16+3=49/16 64^4=16777216 (49/16)/(16777216)=0.0000001825392246246337890625

OpenStudy (anonymous):

have you mistake in your question?

OpenStudy (anonymous):

my second answer x ^3=10 ^4/8

OpenStudy (anonymous):

x^3=(10^4)/8 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

10^4=10000 10^4/8=1250 x=10.772173450159418608796467832597

OpenStudy (anonymous):

if your question was x^3=(10^3)/8 the answer has been x=15 x^3=125 x=5